What is the applied force needed on 2.0 kg block of wood accelerating at 4.0m/s^2 along rough table with a 10 N force of friction acting in the opposite direction?
1
Expert's answer
2021-10-24T18:27:28-0400

Apply Newton's second law to get the answer!


"ma=F-f,\\
F=ma+f=2·4+10=18\text{ N}."